# KineDepth: Utilizing Robot Kinematics for Online Metric Depth Estimation

### v1

@@ III METHODS
This section outlines the pipeline for Monocular Metric Depth Estimation (MMDE) as shown in Fig. 2. At the core of this method is updating a depth regressor in an online manner which maps relative depth to metric depth using the visual measurements of the robot manipulator with known kinematics as known geometry.
@@ III-A Depth Regressor
The depth regressor, $Z(u,v)=f(r(u,v),\bm{\beta})$ for pixel location $[u,v]^{\top}\in\mathbb{R}^{2}$, in our system is a function, $f(\cdot)$, that converts the relative depth image, $r(\cdot)$, obtained from a monocular depth estimation model, such as Depth Anything v2 model [21], into metric depth, $Z(\cdot)$. The depth regressor is parameterized by $\beta$ which we estimate in an online fashion using the robot as a geometric reference. We constrained the depth regressor, $f(\cdot)$, to simple models to ensure stability when applying the regressor to downstream control tasks and evaluated various functions, including power law, rational, Gaussian, logarithmic, linear, and polynomial regressors. We ran an experiment on fitting Depth Anythingv2 predictions to ground truth using data colelcted from our experiments and SciPy’s optimize.fit function. The polynomial regressor demonstrated the highest accuracy in fitting as shown in Tab. I.
@@
TABLE I: Table shows accuracy fit for different regressor to fit the relative depth to estimate metric depth
@@ III-B3 Hydrid
2:     $\beta_{0}\leftarrow[0,1,0],\textbf{P}_{0}\leftarrow\Sigma_{0}$triangleright Kalman Estmation
@@ III-C End Effector Control
where $\gamma^{x}_{t}$ is the end-effector keypoint. Finally, we define the action space as change in end-effector, i.e. $\textbf{x}_{t}=\textbf{x}_{t-1}+\textbf{u}_{t}$. A Linear Quadratic Regulator (LQR) is solved at every timestep with the following cost
@@
where $\mathbf{Q}$ and $\mathbf{R}$ are weighting matrices. Only one action is taken after solving the LQR and then the entire process is repeated (compute goal, end-effector, and LQR) after estimating the latest depth regressor parameters, $\beta_{t}$.
Figure 4: Physical Experiments: The upper row displays successful pick-and-place executions, while the bottom row illustrates failure cases. The figure covers a range of scene configurations and varying levels of grasping difficulty. Failures are primarily due to incorrect depth estimation, or target poses prediction errors, leading to either missed grasps or inaccurate positioning of the marker.
@@ IV-A Implementation details
In the LSTM implementation, we configured the key hyperparameters as follows. The learning rate was set to $1\times 10^{-2}$, with a batch size equal to the number of visible keypoints $M$ in the current frame. The hidden state size was fixed at 128, and the network was trained for $\tau$ 20 time steps per update. $\Sigma_{0}$ is a identity matrix, $\Sigma_{t}$ are a diagonal matrix with 0.5 and $\Sigma_{o}$ are a diagonal matrix with 0.03. In Huber loss, the threshold is set at $\delta=1.0$, while $\alpha_{1}$ and $\alpha_{2}$ are set to 1.0. Meanwhile, $N$ is set to 10 to store the trajectory history for the input of LSTM.
@@ V Discussion and Conclusion
In conclusion, the proposed KineDepth framework introduces a depth regressor that converts relative depth $r$ to the metric depth with high accuracy in the task space region. Further, using Kalman filtering to smooth out the depth regressor predicted from the LSTM network boosted our performance in getting more accurate estimation. KineDepth delivers improved metric depth accuracy in the task space region and it achieved 1.5cm less error for the overall task space region as compared to Metric 3D v2. KineDepth (hybrid) provides 22.1% higher accuracy than the state-of-the-art MMDE, achieving the lowest overall scene error of 0.053 m across all key points.

### v2

@@ III METHODS
This section outlines the pipeline for Monocular Metric Depth Estimation (MMDE) as shown in Fig. 2. At the core of this method is updating a depth regressor in an online manner, which maps relative depth to metric depth using the visual measurements of the robot manipulator with known kinematics as known geometry.
@@ III-A Depth Regressor
The depth regressor, $Z(u,v)=f(r(u,v),\bm{\beta})$ for pixel location $[u,v]^{\top}\in\mathbb{R}^{2}$, in our system is a function, $f(\cdot)$, that converts the relative depth image, $r(\cdot)$, obtained from a monocular depth estimation model, such as Depth Anything v2 model [21], into metric depth, $Z(\cdot)$. The depth regressor is parameterized by $\beta$ which we estimate in an online fashion using the robot as a geometric reference. We constrained the depth regressor, $f(\cdot)$, to simple models to ensure stability when applying the regressor to downstream control tasks and evaluated various functions, including power law, rational, Gaussian, logarithmic, linear, and polynomial regressors. We ran an experiment on fitting Depth Anythingv2 predictions to ground truth using data collected from our experiments and SciPy’s optimize.fit function. The polynomial regressor demonstrated the highest accuracy in fitting, as shown in Tab. I.
@@
TABLE I: Table shows the accuracy fit for different regressor to fit the relative depth to estimate the metric depth
@@ III-B3 Hybrid
2:  $\beta_{0}\leftarrow[0,1,0],\textbf{P}_{0}\leftarrow\Sigma_{0}$triangleright Kalman Estimation
@@ III-C End Effector Control
where $\gamma^{x}_{t}$ is the end-effector keypoint. Finally, we define the action space as a change in end-effector, i.e., $\textbf{x}_{t}=\textbf{x}_{t-1}+\textbf{u}_{t}$. A Linear Quadratic Regulator (LQR) is solved at every timestep with the following cost
@@
where $\mathbf{Q}$ and $\mathbf{R}$ are weighting matrices. Only one action is taken after solving the LQR, and then the entire process is repeated (compute goal, end-effector, and LQR) after estimating the latest depth regressor parameters, $\beta_{t}$.
Figure 4: Physical Experiments: The upper row displays successful pick-and-place executions, while the bottom row illustrates failure cases. The figure covers a range of scene configurations and varying levels of grasping difficulty. Failures are primarily due to incorrect depth estimation or target poses prediction errors, leading to either missed grasps or inaccurate positioning of the marker.
@@ IV-A Implementation details
In the LSTM implementation, we configured the key hyperparameters as follows. The learning rate was set to $1\times 10^{-2}$, with a batch size equal to the number of visible keypoints $M$ in the current frame. The hidden state size was fixed at 128, and the network was trained for $\tau$ 20 time steps per update. $\Sigma_{0}$ is a identity matrix, $\Sigma_{t}$ are a diagonal matrix with 0.5 and $\Sigma_{o}$ are a diagonal matrix with 0.03. In Huber loss, the threshold is set at $\delta=1.0$, while $\alpha_{1}$ and $\alpha_{2}$ are set to 1.0. Meanwhile, $N$ is set to 10 to store the trajectory history for the input of the LSTM.
@@ V Discussion and Conclusion
In conclusion, the proposed KineDepth framework introduces a depth regressor that converts relative depth $r$ to the metric depth with high accuracy in the task space region. Further, using Kalman filtering to smooth out the depth regressor predicted from the LSTM network boosted our performance in getting a more accurate estimation. KineDepth delivers improved metric depth accuracy in the task space region, and it achieved 1.5cm less error for the overall task space region as compared to Metric 3D v2. KineDepth (hybrid) provides 22.1% higher accuracy than the state-of-the-art MMDE, achieving the lowest overall scene error of 0.053 m across all key points.